OpenStudy (anonymous):

im pulling my hair out here in rochester ny the balanced equation is N2 (2 is lower) + O2 (2 is lower) ARROW 2 NO (That is the equation) ok.....Now how many grams of NO are formed from 22.8g of N2????

OpenStudy (aaronq):

Okay, the general process for any stoichiometric problem follows this pattern: 1. Balance the equation 2. Convert your reactants to moles 3. Write a ratio that relates the moles of what you're interested in and their coefficients ---------------------------------------------------------------- e.g. for a general reaction: \(\sf \large \color{red}{a}A + \color{blue}{b}B\rightleftharpoons \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\sf \dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\sf\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) ---------------------------------------------------------------- 4. convert moles of product (or whatever you found) to whatever they're asking for, e.g. mass, volume, etc.

OpenStudy (anonymous):

thank you so much!!!!!!

OpenStudy (aaronq):

no problem !